There are 20 girls and 15 boys in a class.  What is the ratio of a number of girls to the number of boys?

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Given
Number of girls = 20 girls
Number of boys = 15 boys
Total number of students = 20 + 15
= 35
Ratio of number of girls to number of boys = 20 / 15 = 4 / 3

Thanks
Shubham dangi
3 years ago
Hello dear, 
Given that -
No. Of boys =15
No . Of girls=20
Rotio of girls over boys =20/15=4/3
So it's your answer
I hope that you find it good